### v2

# Multimodal Industrial Anomaly Detection by Crossmodal Reverse Distillation

Abstract
Knowledge distillation (KD) has been widely studied in unsupervised image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information. In this paper, we propose Crossmodal Reverse Distillation (CRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code will be available.
1 Introduction
Industrial Anomaly Detection (AD) is an important task aimed at ensuring product quality by identifying defects in products. In real-world industrial scenarios, the majority of object samples are normal, and anomalous samples are often difficult to obtain. Hence, unsupervised AD, which relies only on normal samples for training, has become one of the key research directions in industrial applications [6, 8, 39]. Although traditional RGB-based analysis methods meet the industrial anomaly detection needs to a certain extent, they often struggle to effectively identify anomalies in complex industrial environments, such as in cases of large changes in illumination, or surface bumps that are indistinguishable from color differences. Given the limitations of RGB anomaly detection, data from 3D sensors, including depth map and point cloud, have been widely applied in AD in recent years, which results in growing research interest in unsupervised multimodal AD.
Figure 1: Qualitative results on MVTec 3D-AD [7]. From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [26] with more false negatives, Crossmodal Feature Mapping [12] (abbreviated as CFM) with more false positives, and our method CRD with superior segmentation.
Figure 2: Comparison of our proposed CRD with other multimodal AD methods based on feature learning.
Early unsupervised multimodal AD methods often relies on memory banks, such as BTF [20] and M3DM [31]. These methods construct large memory banks by processing and storing multimodal features of normal samples. During inference, extracted features are compared with normal features in the memory bank to determine if these samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
Recent studies introduce the idea of feature learning to multimodal AD to avoid large-scale storage [26, 15, 12]. The feature learning methods typically leverage pretrained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks. Since only normal samples are used during training, the trainable networks only learn the normality representation ability from the frozen networks. In other words, the abnormality representation ability is not transferred. During inference, the distances between features from the frozen and trainable networks are used to detect anomalies.
Currently, feature learning-based methods are divided into two categories as in Fig. 2: (1) Methods like AST [26] and MMRD [15] (Fig. 2 (a)) follow Knowledge Distillation, where student networks learn the multimodal feature representation ability from teacher networks on normal samples. However, these methods may smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies and leading to false negatives, particularly when one modality is normal and the other is anomalous. (2) Crossmodal Feature Mapping [12] (Fig. 2 (b)) introduces a new paradigm, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one. Each modality has its own distillation target, ensuring the sensitivity to anomalies within each modality. Additionally, to tackle the underfitting problem of Crossmodal Feature Mapping, we propose not to directly use the feature similarity from crossmodal mapping for anomaly map generation. Instead, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality.
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [13], named Crossmodal Reverse Distillation (CRD). CRD naturally combines the strengths of both feature learning paradigms. First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two crossmodal assistants within each modality branch. One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier, which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches. Experimental results on representative multimodal AD datasets demonstrate the effectiveness of CRD in multimodal anomaly detection and localization.
2 Related Work
Unsupervised Anomaly Detection. Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [5, 1, 38, 32, 19, 18]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [25, 2, 22], comparing test samples with stored normal features to detect anomalies. Recently, significant progress has been made in the research of synthetic anomaly images [21, 34, 23], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [27, 30, 26, 3, 13, 28, 16, 14, 37, 17] based on the teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
Multimodal Anomaly Detection. With the release of multiple 3D industrial anomaly detection datasets [7, 9, 24], unsupervised multimodal AD has gradually become a topic of research. Some methods [11, 29], such as BTF [20] and M3DM [31], follow PatchCore [25] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [10, 35, 36] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [4], AST [26], and MMRD [15], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [12] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
3 Preliminaries
Knowledge distillation (KD) is a widely recognized and researched paradigm for unsupervised industrial image AD, typically based on a teacher-student network framework. The teacher network is usually a pretrained model, while the student network is a trainable network identical or similar to the teacher network. During training, only normal samples are used, and the teacher network’s ability to represent normal features is transferred to the student network. This process is often realized using cosine similarity. Let $F_{T}$ represent the feature output of the teacher network, and $F_{S}$ represent the feature output of the student network. The optimization loss is generally expressed as
$\mathrm{Sim}(f_{1},f_{2})=\frac{f_{1}\cdot f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
$\mathcal{L}=\sum_{i=1}^{l}\{1-\mathrm{Sim}(F_{T}^{i},F_{S}^{i})\}$ (2)
where $l$ represents the number of selected feature layers, typically set to 3. During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map. , and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample. When the input is a normal sample, the student’s features are similar to the teacher’s, thus both the values in $M$ and $S$ are low. When the input is an anomalous sample, since the student has not learned the teacher’s abnormal feature representation ability, the feature distance $M$ is large in the anomalous regions, allowing for the detection of both the presence and location of anomalies.
Reverse distillation (RD) [13] is an advanced KD-based AD method. Unlike traditional forward distillation, RD uses an encoder-decoder architecture. The teacher network is a pre-trained encoder WideResNet-50 [33], while the student network is a trainable decoder that generates features matching the size of the teacher’s features. RD also includes a trainable one-class bottleneck embedding (OCBE) module to compress the teacher’s features brefore passing them to the student. During training, the student network learns to reconstruct the teacher’s normal features using normal samples. During inference, the teacher network, with its strong feature extraction ability, captures abnormal patterns in the anomalous input and generates anomaly features. At the same time, the OCBE module compresses the teacher features, preventing abnormal information from inputting student network. Hence, even for anomalous samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
4 Crossmodal Reverse Distillation
Figure 3: Pipeline of our proposed method. The feature evolution visualization highlights that anomalies in each modality are effectively captured. In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
Figure 4: Information flow of the branch of Modality A, showing how Crossmodal Filter and Crossmodal Amplifier incorporate information from Modality B. Crossmodal Filter aids the student decoder in generating normal features by utilizing features from Modality B, while Crossmodal Amplifier strengthens the perception of anomalies in Modality A by amplifying anomaly signals from Modality B.
In the multimodal AD task, each input consists of multiple modalities of data. In this section, we primarily focus on 2D (e.g., RGB images) and 3D (e.g., depth maps) modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{2D},\mathcal{I}^{3D}\}$, where $\mathcal{I}^{2D}$ and $\mathcal{I}^{3D}$ are the 2D and 3D modalities respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which includes both normal and anomalous samples.
4.1 Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated knowledge distillation, which are broadly classified into two categories: (1) Some methods add or concatenate multimodal data together at the input stage, transforming the multi-modal into a single-modal problem [26]. (2) Others employ multiple teacher networks to independently extract features from different modalities, and then merges the features as the learning target for the distillation of student network [15]. The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features from the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
However, the effectiveness of these methods relies on a key assumption: the fused teacher features exhibit strong sensitivity to anomalies, which means they generate a significant difference between normal and anomalous regions. The issue is that in multimodal AD, it is common for one modality to be anomalous while another modality remains normal. In such cases, if the normal modality dominates the fusion, the abnormal information becomes smoothed. As a result, the fused teacher features may fail to capture the abnormal characteristics, remaining close to the normal representation. This causes the student network to fit the teacher’s features even in the anomalous regions, leading to false negatives in anomaly detection.
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 3, which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
For the design of each branch, we draw inspiration from RD. Both 2D and 3D branches consist of four components: the frozen pre-trained teacher encoder, the trainable student decoder, and two crossmodal assistants including Crossmodal Filter and Crossmodal Amplifier (Sec. 4.2). The teacher encoders in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet. The student decoders follow RD, using a structure symmetric to that of the teacher networks. Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to optimize the output features, enabling crossmodal mapping and information fusion between modalities. The anomaly map for each branch is obtained by calculating the cosine distance between the output features of Crossmodal Amplifier and the teacher encoder.
4.2 Crossmodal Assistants
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two crossmodal assistants serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
Crossmodal Filter. The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
Fig. 4 depicts the structure and information flow of Crossmodal Filter (discussed here with modality B assisting modality A as an example). Our proposed Crossmodal Filter module consists of two parts: Bottleneck Projection (BP) and a modified version of the OCBE module. First, BP module aligns the teacher encoder features of modality B to the teacher features of modality A, preventing interference in the decoder process of modality A during later feature fusion. Then, the aligned modality B features are concatenated with the original teacher features of modality A. The OCBE module from RD is modified to retain its feature compression capability while expanding the input feature dimensions to accommodate the fused features.
It is worth noting that to ensure the features of modality B used for assistance are normal, we mimic the concept of OCBE and design Bottleneck Projection as a two-step process of compression and restoration. First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $8\times 8$ by convolution. Then, these features are upsampled back to their original size using deconvolution to obtain ${F^{B}_{BP}}^{i},i=\{1,2,3\}$. During training, we minimize the distance between $F^{B}_{BP}$ and the teacher encoder features of modality A $F^{A}_{E}$, mapping modality B’s features into modality A’s feature space. During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
Crossmodal Amplifier. When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Sec. 4.1. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
As illustrated in Fig. 4, we design Crossmodal Amplifier. First, Inverted Bottleneck Projection (IBP) is designed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$. Then, the mapped features are then dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by
${F^{A}_{CA}}^{i}=\frac{\exp({w_{1}^{A}}^{i})*{F^{A}_{D}}^{i}+\exp({w_{2}^{A}}^{i})*{F^{B}_{IBP}}^{i}}{\exp({w_{1}^{A}}^{i})+\exp({w_{2}^{A}}^{i})}$ (3)
where $w_{1}^{A}$ and $w_{2}^{A}$ are trainable parameters, initially set to 1.0 during training.
Inverted Bottleneck Projection only learns to map modality B’s normal features to modality A’s feature space during training. It means that IBP is not able to make the correct feature mapping when modality B has anomalies in the test sample. Therefore, the fused feature ${F^{A}_{CA}}$ includes modality B’s anomalous information, which leads to a larger distance between the teacher encoder’s features $F^{A}_{E}$ and ${F^{A}_{CA}}$, allowing the anomaly in modality B to be detected in modality A’s branch.
In contrast to Bottleneck Projection in CF, the proposed Inverted Bottleneck Projection first expands the dimensions of features and then compresses them to the original sizes. This process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information during feature mapping.
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Avg.
DepthGAN [7] 53.8/42.1 37.2/42.2 58.0/77.8 60.3/69.6 43.0/49.4 53.4/25.2 64.2/28.5 60.1/36.2 44.3/40.2 57.7/63.1 53.2/47.4
DepthAE [7] 64.8/43.2 50.2/15.8 65.0/80.8 48.8/49.1 80.5/84.1 52.2/40.6 71.2/26.2 52.9/21.6 54.0/71.6 55.2/47.8 59.5/48.1
DepthVM [7] 51.3/38.8 55.1/32.1 47.7/19.4 58.1/57.0 61.7/40.8 71.6/28.2 45.0/24.4 42.1/34.9 59.8/26.8 62.3/33.1 55.5/33.5
BTF [20] 91.8/97.6 74.8/96.9 96.7/97.9 88.3/97.3 93.2/93.3 58.2/88.8 89.6/97.5 91.2/98.1 92.1/95.0 88.6/97.1 86.5/95.9
M3DM [31] 99.4/97.0 90.9/97.1 97.2/97.9 97.6/95.0 96.0/94.1 94.2/93.2 97.3/97.7 89.9/97.1 97.2/97.1 85.0/97.5 94.5/96.4
Shape-guided [11] 98.6/98.1 89.4/97.3 98.3/98.2 99.1/97.1 97.6/96.2 85.7/97.8 99.0/98.1 96.5/98.3 96.0/97.4 86.9/97.5 94.7/97.6
AST [26] 98.3/97.0 87.3/94.7 97.6/98.1 97.1/93.9 93.2/91.3 88.5/90.6 97.4/97.9 98.1/98.2 100/88.9 79.7/94.0 93.7/94.4
MMRD [15] 99.9/98.6 94.3/99.0 96.4/99.1 94.3/95.1 99.2/99.0 91.2/90.1 94.9/99.0 90.1/99.0 99.4/98.7 90.1/98.2 95.0/97.6
CFM [12] 98.8/98.0 87.5/96.6 98.4/98.2 99.2/94.7 99.7/95.9 92.4/96.7 96.4/98.2 94.9/98.3 97.9/97.6 95.0/98.2 96.0/97.2
CRD (Ours) 99.6/97.7 96.1/98.1 98.2/99.1 98.5/98.2 99.7/98.7 90.1/91.0 99.2/99.4 97.6/99.5 100/98.2 83.3/97.4 96.2/97.7
Table 1: Anomaly detection and localization results I-AUROC/PRO (%) on MVTec 3D-AD with the best in bold and the second-best underlined.
4.3 Training Objectives
The overall optimization objective of our method is composed of three main components. Firstly, similar to the approach in RD, we optimize the features output by the student decoder using cosine similarity as
$\mathcal{L}^{2D/3D}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{D}^{2D/3D}}^{i})\}$ (4)
where $\mathcal{L}^{2D}_{D}$ and $\mathcal{L}^{3D}_{D}$ are the losses corresponding to 2D and 3D branches respectively.
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ within these filters to the encoder features. The training losses of 2D and 3D branches are calculated as:
$\mathcal{L}^{2D/3D}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{BP}^{3D/2D}}^{i})\}$ (5)
Finally, Crossmodal Amplifier needs to be optimized,, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. The total loss function is
$\mathcal{L}^{2D/3D}_{CA}=\mathcal{L}^{2D/3D}_{IBP}+\mathcal{L}^{2D/3D}_{output}$ (6)
$\mathcal{L}^{2D/3D}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{IBP}^{3D/2D}}^{i})\}$ (7)
$\mathcal{L}^{2D/3D}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i})\}$ (8)
The overall loss $\mathcal{L}_{CRD}$ is expressed as
$\mathcal{L}_{CRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.
4.4 Anomaly Scoring
Following previous KD-based methods, we generate the anomaly map for each branch using cosine distance as
$\mathcal{M}^{2D/3D}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i}))$ (10)
For the fusion of anomaly maps of two branches, we refer to the method in [8]: First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{2D/3D}$ and $\sigma^{2D/3D}$, respectively. Then, during inference, we normalize the anomaly maps for both branches as
$\mathcal{M}^{2D/3D}_{norm}=\frac{\mathcal{M}^{2D/3D}-\mu^{2D/3D}}{\sigma^{2D/3D}}$ (11)
Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{2D}_{norm}+\mathcal{M}^{3D}_{norm}$. And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
M3DM [31] 62.4/90.6 95.8/92.3 95.8/80.3 100/98.3 88.6/85.5 75.8/68.8 94.9/88.0 83.6/90.6 100/96.6 100/95.5 89.7/88.2
AST [26] 57.4/51.4 74.7/83.5 74.7/71.4 88.9/90.5 59.6/58.7 61.7/59.0 81.6/73.6 84.1/76.9 98.7/91.8 98.7/87.8 78.0/74.4
MMRD [15] 85.4/97.5 100/97.0 94.6/94.2 99.8/98.5 90.8/91.7 74.7/68.0 96.6/97.0 98.4/94.1 100/99.0 100/99.2 94.0/93.6
CFM [12] 68.0/94.2 93.1/90.2 95.2/83.1 88.0/96.5 86.5/87.5 78.2/76.2 91.7/79.1 84.0/91.3 99.8/93.9 96.2/94.9 88.1/88.7
CRD (Ours) 86.7/98.0 100/94.1 98.1/96.8 98.1/98.1 87.2/93.6 72.8/77.3 97.1/98.1 93.0/94.4 100/98.9 99.5/98.1 93.3/94.7
Table 2: Anomaly detection and localization results I-AUC/PRO (%) on Eyecandies with the best in bold and the second-best underlined.
5 Experiments
5.1 Experimental Settings
Dataset. We conduct experiments MVTec 3D-AD [7] and Eyecandies [9]. MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects. The training set consists of 2,656 normal samples, each including an RGB image and corresponding 3D scan information, while the validation set contains 294 normal samples. The test set comprises 1,197 samples, including 948 anomalous samples and 249 normal samples, with pixel-level annotations provided for evaluation. For our experiments, we use the processed depth maps derived from the 3D scans as the 3D modality. Eyecandies is a synthetic dataset of 10 candy categories in an industrial conveyor scenario, providing RGB images, depth maps, and surface normal maps. It contains 10k training samples, 1k validation samples, and 4k test samples. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
Evaluation Metrics. The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP). For anomaly localization, pixel-level AUC (P-AUC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [7] for anomaly localization, which treats anomaly regions of different size equally.
Implementation Details. All RGB images, depth maps and normal maps are scaled to $256\times 256$. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet-50 [33]. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
Figure 5: Qualitative ablation results of crossmodal assistants on MVTec 3D-AD.
5.2 Main Results
Results on MVTec 3D-AD. Tab. 1 shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method CRD with several state-of-the-art 3D anomaly detection methods. CRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%. Notably, CRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach. A subset of qualitative results on MVTec 3D-AD is provided in Fig. 1, showing that our proposed CRD accurately localizes the anomaly regions.
Results on Eyecandies. We evaluate our method on the Eyecandies dataset, as shown in Tab. 2. Our method CRD significantly outperforms all non-KD-based multimodal anomaly detection methods. While it slightly lags behind MMRD [15] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
Method MVTec 3D-AD (%) Eyecandies (%) Avg. (%) Frame Rate (fps)
I-AUC P-AUC PRO I-AUC P-AUC PRO
AST [26] 93.7 97.6 94.4 78.0 90.2 74.4 88.1 13.9
MMRD* [15] 95.0 99.2 97.6 94.0 98.3 93.6 96.3 -
CFM [12] 96.0 99.3 97.2 88.1 97.4 88.7 94.5 12.0
CRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
Table 3: Comparison of feature learning-based methods on MVTec 3D-AD and Eyecandies datasets under the same hardware setup. * indicates that MMRD has no reproducible code.
Comparison with Feature Learning-based Methods. We further compare CRD with feature learning-based methods including AST[26], MMRD[15], and CFM [12] on MVTec 3D-AD and Eyecandies. As in Table 3, CRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps. Despite the multi-branch design, CRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
5.3 Ablation Studies
Branch I-AUC P-AUC P-PRO I-AP P-AP Avg.
2D 86.8 98.9 95.4 95.8 37.7 82.9
2D+Crossmodal 88.9 99.0 95.8 96.5 39.4 83.9
3D 87.1 96.2 88.8 95.2 33.7 80.2
3D+Crossmodal 87.3 97.1 90.1 95.7 34.1 80.9
2D+3D 93.9 99.3 97.4 98.0 45.2 86.8
2D+3D+Crossmodal 96.2 99.4 97.7 98.8 45.5 87.5
Table 4: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
CF CA I-AUC P-AUC P-PRO I-AP P-AP Avg.
- - 93.9 99.3 97.4 98.0 45.2 86.8
- 94.2 99.3 97.5 98.2 45.8 87.0
- 95.6 99.4 97.7 98.6 45.5 87.4
96.2 99.4 97.7 98.8 45.5 87.5
Table 5: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
Study on Multi-branch Distillation. We conduct an ablation study in Tab. 4, which demonstrates that the multi-branch setup, combining both 2D and 3D modalities, significantly outperforms individual modality branches. The I-AUC metric of 93.9% for using both branches is higher than 86.8% for 2D branch only and 87.1% for 3D branch only, highlighting the complementary benefits of modality fusion.
Study on Crossmodal Assistants. We perform an ablation study to evaluate the impact of crossmodal assistants on anomaly detection performance, as shown in Tab. 5. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves the I-AUC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
We also provide visualizations of anomaly maps before and after incorporating crossmodal assistants, as in Fig. 5. Without crossmodal assistants, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing crossmodal assistants, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
Downsampling Size I-AUC P-AUC P-PRO I-AP P-AP Avg.
Identity 93.9 99.3 97.4 98.0 45.8 86.9
$16\times 16$ 93.6 99.3 97.4 97.9 45.2 86.7
$8\times 8$ 94.2 99.3 97.5 98.2 45.8 87.0
$4\times 4$ 93.8 99.3 97.4 98.0 45.3 86.8
Table 6: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
Study on Bottlneck Projection in Crossmodal Filter. We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Tab. 6. When Bottleneck Projection reduces the feature size to $8\times 8$, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $4\times 4$) leads to information loss, while small downsampling size (e.g., $16\times 16$) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
Channel Expansion I-AUC P-AUC P-PRO I-AP P-AP Avg.
Identity 95.7 99.3 97.7 98.7 45.1 87.3
2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5 87.4
4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0 87.3
Table 7: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
Study on Inverted Bottlneck Projection in Crossmodal Amplifier. As recorded in Tab. 7, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier. The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency. This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
Aggregation I-AUC P-AUC P-PRO I-AP P-AP Avg.
$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5 87.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4 87.1
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 94.3 99.3 97.2 98.2 43.5 86.5
Table 8: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
Study on Anomaly Scoring. Tab. 8 demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps. Normalization ensures that anomaly maps from different modalities are on the same scale, making fusion more effective.
Figure 6: Visualization of the features output by CF and CA.
Visualization of Mapped Features. We visualizes the output features of CF and CA using the visualization decoders [32] in Fig. 6, which intuitively demonstrates the roles of CA and CF. Regardless of whether the initial modality is normal or anomalous, CF consistently produces normal outputs, enabling the student decoder of the other modality to generate anomaly-free results. When the initial modality is anomalous, CA introduces discrepancies with the other modality, amplifying anomalies for enhanced detection.
6 Conclusion
In this paper, Crossmodal Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal AD task. CRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch. Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities. Experimental results show that CRD achieves SOTA performance in both anomaly detection and localization.
Supplementary Material
Overview
This supplementary material provides extended analyses and visualizations to complement the main paper:
  • Detailed P-AUC results for anomaly localization on MVTec 3D-AD and Eyecandies datasets. (Appendix A)
  • Training accuracy curves to demonstrate the model’s convergence over training. (Appendix B)
  • Detailed results of ablation studies on MVTec 3D-AD dataset. (Appendix C)
  • Additional qualitative visualizations on MVTec 3D-AD and Eyecandies datasets. (Appendix D)
Appendix A P-AUC Metric
A.1 P-AUC on MVTec 3D-AD
Method/Category Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
BTF [20] 99.6 99.2 99.7 99.4 98.1 97.4 99.6 99.8 99.4 99.5 99.2
M3DM [31] 99.5 99.3 99.7 98.5 98.5 98.4 99.6 99.4 99.7 99.6 99.2
AST [26] - - - - - - - - - - 97.6
MMRD [15] - - - - - - - - - - 99.2
CFM [12] 99.7 99.2 99.9 97.2 98.7 99.3 99.8 99.9 99.8 99.8 99.3
CRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
Table 9: Pixel-level anomaly localization results P-AUC (%) on MVTec 3D-AD with the best in bold.
This section compares our method with other mainstream multimodal AD methods on MVTec 3D-AD dataset using the P-AUC metric, including memory bank-based methods BTF [20] and M3DM [31], KD-based methods AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. As shown in Tab. 9, our method achieves SOTA performance, reaching 99.4% on the P-AUC metric. Notably, for certain categories like Dowel, our CRD significantly outperforms previous methods. In most other categories, it consistently maintains competitive or near-optimal performance.
A.2 P-AUC on Eyecandies
Methods Candy Cane Chocolate Cookie Chocolate Praline Confetto Gummy Bear Hazelnut Truffle Licorice Sandwich Lollipop Marsh- mallow Peppermint Candy Avg.
RGB-D [9] 97.3 92.7 95.8 94.5 92.9 80.6 82.7 97.7 93.1 92.8 92.0
RGB-cD-n [9] 98.0 97.9 98.2 97.8 95.1 85.3 97.1 97.8 98.5 96.7 96.2
M3DM [31] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
AST [26] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
MMRD [15] - - - - - - - - - - 98.3
CFM [12] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
CRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
Table 10: Pixel-level anomaly localization results P-AUC (%) on Eyecandies with the best in bold.
This section presents the quantitative results of anomaly localization on Eyecandies dataset using the P-AUC metric in Tab. 10. The primary comparison methods include RGB-D [9] and RGB-cD-n [9], proposed alongside the dataset, memory bank-based methods such as BTF [20] and M3DM [31], KD-based methods like AST [26] and MMRD [15], and Crossmodal Feature Mapping CFM [12]. Our method significantly outperforms all non-KD-based multimodal anomaly detection methods, achieving 98.5%.
Appendix B Training Curves: I-AUC, P-AUC, and PRO
(a) I-AUC.
(b) P-AUC.
(c) PRO.
Figure 7: AD performance over 500 training epochs.
We provide training curves of I-AUC, P-AUC, and PRO for the carrot and peach categories from the MVTec 3D-AD dataset. As shown in Fig. 7, all metrics stabilize after approximately 50 epochs, indicating that the model converges efficiently without significant performance fluctuations. This case demonstrates the robustness of our method, as it avoids over-generalization even with extended training. Following the setting of Reverse Distillation [13], we adopt 200 epochs for training to ensure sufficient convergence while maintaining computational efficiency.
Appendix C Additional Results of Ablation Studies
In the main text, we present five aspects of ablation studies:
  1. 1.
    Study on Multi-branch Distillation demonstrates that combining the distillation results from multiple branches of different modalities achieves better performance, highlighting the importance of multiple branches. (Tab. 11)
  2. 2.
    Study on Crossmodal Assistants shows that incorporating crossmodal assistants into Multi-branch Distillation based on Reverse Distillation significantly improves anomaly detection, proving their effectiveness. (Tab. 12)
  3. 3.
    Study on Bottleneck Projection in Crossmodal Filter explores the impact of downsampling sizes in the Bottleneck Projection module of Crossmodal Filter, concluding that a size of $8\times 8$ yields optimal results. (Tab. 13)
  4. 4.
    Study on Inverted Bottleneck Projection in Crossmodal Amplifier investigates the effect of channel expansion in the Inverted Bottleneck Projection module of Crossmodal Amplifier, finding that a 2×2$\times$2 × expansion achieves the best performance. (Tab. 14)
  5. 5.
    Study on Anomaly Scoring examines different methods for fusing anomaly maps from multiple branches, confirming that our selected approach delivers the best results. (Tab. 15)
In the main text, we provide average results across all categories for comparison. Here, we present detailed results for each category, as shown in Tab. 11, Tab. 12, Tab. 13, Tab. 14, and Tab. 15.
Metrics Branch Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC 2D 97.8 94.0 93.9 58.6 98.5 85.3 91.8 62.1 99.2 86.3 86.8
2D+Crossmodal 98.7 97.6 94.2 65.7 98.4 88.1 91.6 66.5 97.9 90.3 88.9
3D 99.1 63.8 97.8 99.5 85.3 73.5 99.4 98.8 98.3 55.2 87.1
3D+Crossmodal 98.3 67.6 98.1 99.1 86.1 76.6 99.2 99.4 98.1 50.5 87.3
2D+3D 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
2D+3D+Crossmodal 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUC 2D 99.3 99.5 99.5 98.0 99.5 96.1 99.4 99.4 99.4 99.1 98.9
2D+Crossmodal 99.4 99.6 99.5 98.5 99.5 95.9 99.4 99.4 99.4 99.2 99.0
3D 99.0 94.0 99.8 99.0 98.7 78.5 99.8 99.9 99.5 93.7 96.2
3D+Crossmodal 98.9 93.8 99.8 98.9 99.2 84.8 99.8 99.9 99.5 95.9 97.1
2D+3D 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
2D+3D+Crossmodal 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO 2D 97.2 98.6 98.3 89.4 97.3 87.3 97.8 97.6 95.5 95.3 95.4
2D+Crossmodal 97.8 98.9 98.2 92.1 97.3 86.6 97.7 97.7 95.5 95.7 95.8
3D 97.0 81.3 99.1 97.6 95.4 47.5 99.3 99.6 96.2 75.3 88.8
3D+Crossmodal 96.7 80.0 99.1 97.5 96.9 51.8 99.4 99.6 96.8 82.9 90.1
2D+3D 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
2D+3D+Crossmodal 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
I-AP 2D 99.4 98.5 98.6 86.9 99.7 96.0 97.8 86.0 99.7 95.3 95.8
2D+Crossmodal 99.7 99.4 98.7 88.3 99.7 97.0 97.7 88.7 99.2 97.0 96.5
3D 99.8 88.9 99.5 99.9 95.7 89.8 99.8 99.7 99.3 79.4 95.2
3D+Crossmodal 99.6 91.2 99.6 99.8 95.7 92.9 99.8 99.9 99.2 79.0 95.7
2D+3D 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
2D+3D+Crossmodal 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
P-AP 2D 49.2 42.9 31.8 47.6 43.6 23.0 37.8 17.9 52.1 31.4 37.7
2D+Crossmodal 52.2 44.8 30.8 52.3 49.9 23.0 35.2 21.0 50.5 34.0 39.4
3D 23.9 8.4 45.7 56.4 21.5 8.9 59.7 59.7 49.2 3.3 33.7
3D+Crossmodal 20.7 2.6 47.3 54.7 26.8 9.9 64.7 58.5 51.5 4.1 34.1
2D+3D 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
2D+3D+Crossmodal 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
Table 11: Ablation study results (%) of Multi-branch Distillation on MVTec 3D-AD.
Metrics CF CA Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC - - 99.7 89.8 97.8 97.1 99.9 81.3 99.4 96.8 99.9 77.2 93.9
- 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
- 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUC - - 99.4 99.3 99.7 99.4 99.6 97.2 99.8 99.9 99.6 99.1 99.3
- 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
- 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO - - 98.2 98.0 99.0 98.2 98.4 90.2 99.3 99.5 98.0 95.4 97.4
- 98.2 98.0 99.0 98.0 98.4 90.5 99.3 99.5 97.9 95.8 97.5
- 97.8 98.3 99.1 98.2 98.6 91.4 99.3 99.5 98.2 96.7 97.7
97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
I-AP - - 99.9 97.3 99.5 99.2 100 93.9 99.9 98.9 99.9 91.3 98.0
- 99.9 97.8 99.5 98.7 99.9 95.5 99.8 99.0 99.9 91.9 98.2
- 99.9 98.5 99.7 99.5 99.9 96.9 99.9 99.2 100 92.6 98.6
99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
P-AP - - 42.3 38.2 41.6 71.4 40.7 26.9 58.8 56.7 53.5 22.2 45.2
- 47.4 37.1 41.4 69.6 41.4 27.8 61.2 57.9 52.2 22.4 45.8
- 30.5 39.4 42.6 69.6 48.8 29.1 61.3 57.4 55.0 20.9 45.5
27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
Table 12: Ablation study results (%) of Crossmodal Assistants on MVTec 3D-AD.
Metrics Downsampling Size Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC Identity 99.6 88.8 97.9 95.6 99.7 84.6 98.8 96.8 99.8 77.1 93.9
$16\times 16$ 99.7 88.1 97.6 96.9 100 81.6 99.2 96.7 99.8 76.7 93.6
$8\times 8$ 99.5 91.3 97.7 95.2 99.7 84.4 99.2 96.6 99.8 78.2 94.2
$4\times 4$ 99.7 89.5 98.1 95.6 99.9 81.6 98.8 96.8 99.8 77.5 93.8
P-AUC Identity 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.1 99.3
$16\times 16$ 99.4 99.3 99.7 99.3 99.7 97.3 99.8 99.9 99.6 99.1 99.3
$8\times 8$ 99.4 99.3 99.7 99.3 99.7 97.2 99.8 99.9 99.6 99.2 99.3
$4\times 4$ 99.4 99.3 99.7 99.3 99.6 97.0 99.8 99.9 99.6 99.1 99.3
PRO Identity 98.2 98.0 99.0 98.1 98.4 90.2 99.3 99.5 98.0 95.6 97.4
$16\times 16$ 98.1 98.0 99.0 98.1 98.4 90.6 99.2 99.5 98.0 95.5 97.4
$8\times 8$ 98.2 98.0 99.0 98.0 98.4 90.5 99.3 99.5 97.9 95.8 97.5
$4\times 4$ 98.1 98.1 99.0 98.0 98.4 89.8 99.2 99.5 98.0 95.7 97.4
I-AP Identity 99.9 96.9 99.5 98.8 99.9 95.5 99.7 99.0 99.9 91.3 98.0
$16\times 16$ 99.9 96.7 99.5 99.2 100 93.9 99.8 98.9 99.9 90.9 97.9
$8\times 8$ 99.9 97.8 99.5 98.7 99.9 95.5 99.8 99.0 99.9 91.9 98.2
$4\times 4$ 99.9 97.3 99.6 98.8 100 93.8 99.7 98.9 99.9 91.8 98.0
P-AP Identity 45.7 37.5 41.2 69.3 40.6 27.8 62.2 57.5 52.8 23.2 45.8
$16\times 16$ 42.7 37.1 41.4 70.2 41.3 27.3 59.6 56.8 52.9 22.3 45.2
$8\times 8$ 47.4 37.1 41.4 69.6 41.4 27.8 61.2 57.9 52.2 22.4 45.8
$4\times 4$ 44.4 38.0 41.6 70.1 40.6 26.4 59.3 56.1 54.0 22.8 45.3
Table 13: Ablation study results (%) of different downsampling sizes in Bottlneck Projection on MVTec 3D-AD.
Metrics Channel Expansion Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC Identity 99.6 95.2 98.3 98.0 99.7 87.4 99.2 97.6 100 81.9 95.7
2×2$\times$2 × 99.7 93.5 98.5 98.1 99.7 88.9 99.5 97.3 100 80.4 95.6
4×4$\times$4 × 99.5 94.7 98.0 98.6 99.7 88.8 99.2 97.5 100 80.8 95.7
P-AUC Identity 99.3 99.4 99.7 99.3 99.7 97.4 99.8 99.5 99.7 99.4 99.3
2×2$\times$2 × 99.3 99.3 99.7 99.3 99.7 97.5 99.8 99.9 99.7 99.4 99.4
4×4$\times$4 × 99.3 99.3 99.7 99.3 99.7 97.6 99.8 99.9 99.7 99.4 99.4
PRO Identity 97.6 98.3 99.1 98.2 98.6 90.9 99.3 99.5 98.2 97.0 97.7
2×2$\times$2 × 97.8 98.3 99.1 98.2 98.6 91.4 99.3 99.5 98.2 96.7 97.7
4×4$\times$4 × 97.7 98.2 99.1 98.2 98.6 91.6 99.3 99.5 98.2 96.8 97.7
I-AP Identity 99.9 98.9 99.6 99.5 99.9 96.5 99.8 99.3 100 93.2 98.7
2×2$\times$2 × 99.9 98.5 99.7 99.5 99.9 96.9 99.9 99.2 100 92.6 98.6
4×4$\times$4 × 99.9 98.7 99.5 99.6 99.9 96.9 99.8 99.3 100 92.8 98.6
P-AP Identity 27.9 39.9 42.1 67.6 49.8 28.4 62.0 57.0 55.4 21.3 45.1
2×2$\times$2 × 30.5 39.4 42.6 69.6 48.8 29.1 61.3 57.4 55.0 20.9 45.5
4×4$\times$4 × 29.0 38.4 42.4 68.1 48.8 29.1 60.9 57.3 55.7 20.4 45.0
Table 14: Ablation study results (%) of different channel expansion sizes in Inverted Bottlneck Projection on MVTec 3D-AD.
Metrics Aggregation Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.6 92.9 98.3 98.3 99.3 90.3 98.9 96.7 100 81.6 95.6
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.6 89.8 98.4 98.9 99.8 82.4 99.0 96.4 100 78.8 94.3
P-AUC $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.4 96.8 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.1
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.4 98.9 99.8 99.3 99.7 97.0 99.8 99.9 99.7 99.5 99.3
PRO $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 97.8 96.8 99.1 98.2 98.7 91.1 99.3 99.5 98.2 97.2 97.6
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 98.0 96.9 99.1 98.2 98.6 88.9 99.4 99.5 98.1 95.7 97.2
I-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 99.9 99.1 99.6 99.6 99.9 97.1 99.8 99.3 100 93.7 98.8
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 99.9 98.2 99.6 99.6 99.8 97.2 99.7 99.0 100 92.7 98.6
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 99.9 97.3 99.7 99.7 100 94.1 99.6 98.9 100 92.5 98.2
P-AP $\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 27.9 36.7 42.8 70.1 50.1 29.4 62.9 57.4 54.8 22.7 45.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 31.6 27.0 43.9 70.4 48.8 29.6 60.3 55.9 55.3 21.2 44.4
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [12] 37.8 28.6 42.5 72.1 45.7 17.6 60.5 54.3 54.2 21.4 43.5
Table 15: Ablation study results (%) of anomaly map aggregation strategy on MVTec 3D-AD.
Appendix D Additional Qualitative Results
Figure 8: Qualitative results on MVTec 3D-AD. From top to bottom: 2D images, 3D depth maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
D.1 Qualitative Results on MVTec 3D-AD
This section first provides visualizations of the anomaly maps generated by each branch and the fused results, as shown in Fig. 8. The visualizations reveal two key findings: (1) In the 3D branch, the significant variation in depth values along edges introduces abundant edge information, making it challenging for student features to fit teacher features at edge regions, leading to false positives. But this issue is less pronounced in the 2D branch. (2) Some anomalies are modality-specific. For instance, discoloration is only detectable in RGB images, while the absence of chocolate on cookies is only evident in depth maps. Consequently, anomaly maps from individual branches may fail to fully capture all anomalies within a sample. However, the fused results effectively achieve comprehensive anomaly localization.
Figure 9: Qualitative results on MVTec 3D-AD. From left to right: the examples of Bagel, Cable Cland, Cookie, Dowel, and Potato.
In the main text, we visualize anomaly maps with and without crossmodal assistants, demonstrating their effectiveness in improving anomaly segmentation accuracy by reducing false positives and false negatives. In this section, we provide additional examples to further validate the robustness of our method. As shown in Fig. 9, the inclusion of CF and CA consistently enhances the precision of anomaly localization across various categories, including bagel, cable cland, cookie, dowel, and potato. These results highlight the critical role of crossmodal interactions in refining anomaly detection.
D.2 Qualitative result on Eyecandies
Figure 10: Qualitative results on Eyecandies. From top to bottom: the input 2D images, the corresponding 3D normal maps, the output anomaly maps of 2D branch, the output anomaly maps of 3D branch, the output anomaly maps of CRD, and the ground truth masks.
Fig. 10 visualizes the anomaly maps generated by our proposed CRD on Eyecandies. Our method demonstrates strong sensitivity to both 2D and 3D modalities, enabling comprehensive multimodal anomaly localization.