### v3

# Tuned Reverse Distillation: Enhancing Multimodal Industrial Anomaly Detection with Crossmodal Tuners

Abstract
Knowledge distillation (KD) has been widely studied in unsupervised image Anomaly Detection (AD), but its application to unsupervised multimodal AD remains underexplored. Existing KD-based methods for multimodal AD that use fused multimodal features to obtain teacher representations face challenges. Anomalies that only exist in one modality may not be effectively captured in the fused teacher features, leading to detection failures. Besides, these methods do not fully leverage the rich intra- and inter-modality information that are critical for effective anomaly detection. In this paper, we propose Tuned Reverse Distillation (TRD) based on Multi-branch design to realize Multimodal Industrial AD. By assigning independent branches to each modality, our method enables finer detection of anomalies within each modality. Furthermore, we enhance the interaction between modalities during the distillation process by designing two Crossmodal Tuners including Crossmodal Filter and Amplifier. With the idea of crossmodal mapping, the student network is allowed to better learn normal features while anomalies in all modalities are ensured to be effectively detected. Experimental verifications on multimodal AD datasets demonstrate that our method achieves state-of-the-art performance in multimodal anomaly detection and localization. Code is available at https://github.com/hito2448/TRD.
Index Terms:
Multimodal anomaly detection, knowledge distillation, unsupervised anomaly detection, anomaly localization.
I Introduction
Industrial Anomaly Detection (AD) is an important task aimed at ensuring product quality by identifying defects in products. In real-world industrial scenarios, the majority of object samples are normal, and anomalous samples are often difficult to obtain. Hence, unsupervised AD, which relies only on normal samples for training, has become one of the key research directions in industrial applications [1, 2, 3]. Although traditional RGB-based analysis methods meet the industrial anomaly detection needs to a certain extent, they often struggle to effectively identify anomalies in complex industrial environments, such as in cases of large changes in illumination, or surface bumps that are indistinguishable from color differences. Given the limitations of RGB anomaly detection, data from 3D sensors, including depth map and point cloud, have been widely applied in AD in recent years, which results in growing research interest in unsupervised multimodal AD.
Figure 1: Qualitative comparisons on MVTec 3D-AD [4]. From top to bottom: 2D images, 3D depth maps, the ground truth masks, the output anomaly maps of multimodal AD methods based on feature learning showing AST [5] with more false negatives, Crossmodal Feature Mapping [6] (abbreviated as CFM) with more false positives, and our method TRD with superior segmentation.
Early unsupervised multimodal AD methods often rely on memory banks, such as BTF [7] and M3DM [8]. These methods construct large memory banks by processing and storing multimodal features of normal samples. During inference, extracted features are compared with normal features in the memory bank to determine if these samples are anomalous. However, such methods inevitably require substantial storage, which becomes a bottleneck, especially when dealing with large-scale datasets.
Recent studies introduce the idea of feature learning to multimodal AD to avoid large-scale storage [5, 9, 6]. The feature learning-based methods typically leverage pre-trained networks which is frozen during training to generate rich feature representations and transfer them to trainable networks. Since only normal samples are used during training, the trainable networks only learn the normality representation ability from the frozen networks. In other words, the abnormality representation ability is not transferred. During inference, the distances between features from the frozen and trainable networks are used to detect anomalies.
Figure 2: Comparison of our proposed TRD with previous KD-based methods. Previous KD-based methods tend to suffer from anomaly smoothing when one modality is normal and the other contains anomalies. For example, in this figure, the anomaly on the bagel is barely visible in the 2D modality and resembles normal samples, while it is clearly distinguishable in the 3D modality. However, one of previous KD-based methods AST [5], which fuses the two modalities before distillation, fails to detect the anomaly and results in false negative, as in (a). In contrast, our TRD employs a multi-branch design and Crossmodal Tuners to preserve modality-specific anomaly cues, enabling more accurate anomaly localization as in (b).
Currently, feature learning-based methods for multimodal AD are primarily divided into two categories. One is Crossmodal Feature Mapping [6], which introduces a novel paradigm for addressing multimodal AD, and realizes multimodal AD by mapping features across modalities to capture crossmodal relationships. When anomalies occur, the failure of feature mapping enables anomaly localization. But inevitably, modal discrepancy between 2D and 3D features results in underfitting and causes false positives in the output anomaly maps.
Knowledge distillation-based methods [5, 9] represent another widely accepted category of feature learning-based methods, which extend the typical KD paradigm, originally used in image anomaly detection, to multimodal settings. In such methods, student networks are trained on normal samples to learn multimodal feature representations from teacher networks. Anomalies are then identified by measuring the feature discrepancies between the student and the teacher. To obtain multimodal teacher features as distillation targets, KD-based multimodal AD methods typically adopt one of two fusion strategies. Some fuse 2D and 3D data at the input level before passing it through a pre-trained encoder. Others extract features from separately frozen 2D and 3D encoders and then merge them. However, both fusion strategies smooth out abnormal features during fusion, reducing the teacher’s sensitivity to anomalies, particularly when one modality is normal and the other is anomalous. In such cases, the fused teacher features tend to be dominated by the normal modality, reducing their sensitivity to subtle anomalies and increasing the risk of false negatives, as illustrated in Fig. 2 (a).
To address the anomaly smoothing issue in previous KD-based methods, our intuition is to expand the single-branch distillation into a multi-branch one. Each modality is assigned an independent distillation objective, allowing the student network to learn from the corresponding teacher features. This multi-branch design preserves the anomaly sensitivity of each modality and avoids the anomaly smoothing problem caused by early feature fusion. With this setup, each branch is able to accurately detect anomalies specific to its modality. However, simply performing late fusion by combining anomaly maps from each branch may still weaken anomaly signals, particularly when anomalies only in one modality are subtle. To further mitigate this issue, we integrate the idea of crossmodal mapping into the multi-branch distillation, where features from another modality’s teacher network, along with the inter-modal relationship, help generate features for the student network of a given modality. This enables each branch not only to retain its own anomaly information but also to capture crossmodal cues, ensuring that anomalies present in any modality are effectively reflected in the final output.
In summary, we propose a targeted extension of the widely studied KD-based AD paradigm Reverse Distillation (RD) [10], named Tuned Reverse Distillation (TRD). TRD naturally combines the strengths of existing feature learning paradigms, as shown in Fig. 2 (b). First, we design Multi-branch Distillation that enhances anomaly detection within each modality and refines the modality fusion process to generate more precise anomaly maps. Next, we introduce two Crossmodal Tuners within each modality branch. One is Crossmodal Filter, which uses information from other modalities to help the student decoder generate normal features. And the other is Crossmodal Amplifier, which allows a modality branch to detect anomalies from other modalities. Finally, anomaly scores of test samples are calculated by integrating anomaly maps of all branches. Experimental results on representative multimodal AD datasets demonstrate the effectiveness of TRD in multimodal anomaly detection and localization.
Our contributions are as follows:
  • We propose Tuned Reverse Distillation based on Multi-branch Distillation for multimodal industrial AD, which effectively detects anomalies in all modalities.
  • Two Crossmodal Tuners are introduced, including Crossmodal Filter that helps the student decoder generate normal features, and Crossmodal Amplifier that amplifies the perception of anomalies from other modalities.
  • Experiments show that our proposed TRD achieves state-of-the-art performance on multimodal AD datasets.
II Related Work
II-A Unsupervised Anomaly Detection
Unsupervised AD has gained increasing attention in recent years. Initially, many unsupervised AD methods relied on generative models [11, 12, 13, 14, 15, 16, 17]. These models are trained on normal samples to learn the ability to reconstruct normal data. During inference, reconstruction errors are used to classify input images as normal or anomalous. Other methods introduce memory banks [18, 19, 20], comparing test samples with stored normal features to detect anomalies. Recently, significant progress has been made in the research of synthetic anomaly images [21, 22, 23], which simulate real-world scenarios to assist in unsupervised AD tasks. Additionally, knowledge distillation methods [24, 5, 10, 25, 26, 27, 28, 29] based on the teacher-student framework have also been applied to unsupervised AD. These methods train the student network to learn the feature representations of the teacher network on normal samples, and then use the differences in feature representations between the teacher and student networks on anomalous pixels to locate anomalies. Due to the intuitive and simple nature of KD methods, KD-based AD methods have become an important focus in the field of unsupervised anomaly detection.
II-B Multimodal Anomaly Detection
With the release of multiple 3D industrial anomaly detection datasets [4, 30, 31], unsupervised multimodal AD has gradually become a topic of research. Some methods [32, 33], such as BTF [7] and M3DM [8], follow PatchCore [18] by leveraging memory banks for 3D AD, adding the storage of 3D features to the original PatchCore method. Other methods [34, 35, 36] rely on reconstruction networks to detect anomalies based on the reconstruction results of multimodal data. Knowledge distillation paradigms have also been explored in some methods, such as 3D-ST [37], AST [5], and MMRD [9], with the goal of having the student network mimic the multimodal features output by the teacher network. Crossmodal Feature Mapping [6] proposes a new multimodal solution, using two neural networks to perform crossmodal mapping and locate anomalies based on the mapping results.
III Preliminaries
Knowledge distillation (KD) is a widely recognized and researched paradigm for unsupervised industrial image AD, typically based on a teacher-student network framework. The teacher network is usually a pretrained model, while the student network is a trainable network identical or similar to the teacher network. During training, only normal samples are used, and the teacher network’s ability to represent normal features is transferred to the student network. This process is often realized using cosine similarity. Let $F_{T}$ represent the feature output of the teacher network, and $F_{S}$ represent the feature output of the student network. The optimization loss is generally expressed as
$\mathrm{Sim}(f_{1},f_{2})=\frac{f_{1}\cdot f_{2}}{\|f_{1}\|\|f_{2}\|}$ (1)
$\mathcal{L}=\sum_{i=1}^{l}\{1-\mathrm{Sim}(F_{T}^{i},F_{S}^{i})\}$ (2)
where $l$ represents the number of selected feature layers, typically set to 3. During inference, the multi-layer cosine distance between the teacher and student features is used, with $\mathcal{M}=\sum(1-\mathrm{Sim}(F_{T},F_{S}))$ as the anomaly map. , and $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score for the sample. When the input is a normal sample, the student’s features are similar to the teacher’s, thus both the values in $M$ and $S$ are low. When the input is an anomalous sample, since the student has not learned the teacher’s abnormal feature representation ability, the feature distance $M$ is large in the anomalous regions, allowing for the detection of both the presence and location of anomalies.
Reverse distillation (RD) [10] is an advanced KD-based AD method. Unlike traditional forward distillation, RD uses an encoder-decoder architecture. The teacher network is a pre-trained encoder WideResNet-50 [38], while the student network is a trainable decoder that generates features matching the size of the teacher’s features. RD also includes a trainable one-class bottleneck embedding (OCBE) module to compress the teacher’s features brefore passing them to the student. During training, the student network learns to reconstruct the teacher’s normal features using normal samples. During inference, the teacher network, with its strong feature extraction ability, captures abnormal patterns in the anomalous input and generates anomaly features. At the same time, the OCBE module compresses the teacher features, preventing abnormal information from inputting student network. Hence, even for anomalous samples, the student network generates anomaly-free features, which enables RD to detect and localize anomalies by calculating the cosine distance between the student and teacher features.
IV Tuned Distillation
Figure 3: Information flow of the branch for Modality A during training. This branch consists of a frozen teacher encoder, a trainable student decoder, and two Crossmodal Tuners including Crossmodal Filter and Crossmodal Amplifier. Crossmodal Filter incorporates mapped features from Modality B to guide the student decoder of Modality A in generating anomaly-free features. Crossmodal Amplifier injects anomaly signals from Modality B after the decoder, enhancing the anomaly sensitivity of Modality A.
In the multimodal AD task, each input consists of multiple modalities of data. In this section, we primarily focus on 2D (e.g., RGB images) and 3D (e.g., depth maps) modalities as examples for studying multimodal anomaly detection. Therefore, each input is represented as =$\mathcal{I}=\{\mathcal{I}^{2D},\mathcal{I}^{3D}\}$, where $\mathcal{I}^{2D}$ and $\mathcal{I}^{3D}$ are the 2D and 3D modalities respectively. The training set consists of a collection of $n$ normal samples, denoted as $\{\mathcal{I}_{train}^{1},\mathcal{I}_{train}^{2},...,\mathcal{I}_{train}^{n}\}$. The test set $\{\mathcal{I}_{test}^{1},\mathcal{I}_{test}^{2},...,\mathcal{I}_{test}^{m}\}$ contains $m$ samples, which include both normal and anomalous samples.
IV-A Multi-branch Distillation
In multimodal industrial AD, several methods have incorporated Knowledge Distillation, which are broadly classified into two categories:
  • Some methods add or concatenate multimodal data at the input stage, transforming the multimodal into a singlemodal problem [5].
  • Others employ multiple teacher networks to independently extract features from different modalities, and then merge the features as the learning target for the distillation of the student network [9].
The core idea of these methods is to train the student network to learn the fused normal features from the teacher networks, thereby learning the normal pattern. During inference, when anomalous samples are input, the student network fails to fit the fused abnormal features of the teacher, allowing anomalies to be detected by measuring the feature distance between the student and teacher networks.
However, the effectiveness of these methods relies on a key assumption: the fused teacher features exhibit strong sensitivity to anomalies, which means they generate a significant difference between normal and anomalous regions. The issue is that in multimodal AD, it is common for one modality to be anomalous while another modality remains normal. In such cases, if the normal modality dominates the fusion, the abnormal information becomes smoothed. As a result, the fused teacher features may fail to capture the abnormal characteristics, remaining close to the normal representation. This causes the student network to fit the teacher’s features even in the anomalous regions, leading to false negatives in anomaly detection.
To resolve the problem of anomaly smoothing during modality fusion, we introduce Multi-branch Distillation (MBD) as in Fig. 2 (b), which differs from previous KD-based multimodal AD methods with a multi-branch design. (In this section, we focus on the example case of dual-branch including 2D and 3D branch, with the multi-branch scenario extending similarly.) Within MBD, each branch corresponds to a specific modality. During training, the pre-trained teacher network independently distills knowledge for each modality. At inference, an anomaly map is generated for each modality, and these maps are fused after normalization. By deferring the fusion of modalities to the inference stage, our method MBD allows for more precise integration of multimodal information, effectively preventing the smoothing of anomalies during feature fusion and reducing the risk of false negatives.
For the design of each branch, we draw inspiration from RD. The two branches corresponding to the 2D and 3D modalities are designed symmetrically. Fig. 3 depicts the detailed architecture of one branch (for modality A), which includes four components: a frozen pre-trained teacher encoder, a trainable student decoder, and two Crossmodal Tuners including Crossmodal Filter and Crossmodal Amplifier (Section IV-B). The teacher encoder in both branches share parameters and use a WideResNet-50 pre-trained on ImageNet. The student decoders in both branches follow RD, using a structure symmetric to that of the teacher networks. Crossmodal Filters replace the OCBE module of RD, while Crossmodal Amplifiers are placed after the student decoders to tune the output features, enabling crossmodal mapping and information fusion between modalities.
Figure 4: Illustration of the Crossmodal Filter within the branch of Modality A, which consists of two components including Bottleneck Projection and OCBE. First, the encoder features from the other modality (Modality B) ${F_{E}^{B}}^{i},i=\{1,2,3\}$ are processed by the Bottleneck Projection to suppress anomaly signals and generate compact, normal-like features. These features are then concatenated with the encoder features from Modality A ${F_{E}^{A}}^{i},i=\{1,2,3\}$ and passed into the OCBE module. The resulting fused representation $F_{CF}^{A}$ retains rich multimodal information while remaining anomaly-free, and is used to guide the student decoder in reconstructing anomaly-free features within the Modality A branch.
IV-B Crossmodal Tuners
To fully leverage the complementary information between modalities, we introduce crossmodal interactions within Multi-branch Distillation. To be specific, we design two Crossmodal Tuners serving two key purposes: (1) Crossmodal Filter (CF) helps the student network of one modality reconstruct normal features by utilizing information from another modality, even when that modality exhibits anomalies. (2) Crossmodal Amplifier (CA) allows the abnormal information from the anomalous modality to be reflected in the student features from the normal modality. With CA, the anomaly is more prominent in both branch even when one of the modalities is normal, thus prevented from being overlooked in the final anomaly map.
IV-B1 Crossmodal Filter
The training objective of RD is to ensure that the student decoder generates anomaly-free features. It is achieved through the OCBE bottleneck module in vanilla RD. By further compressing the encoder features, the OCBE module effectively prevents anomaly disturbances from propagating into the student decoder, ensuring that the student decoder generates anomaly-free features even when anomaly samples are input during inference. However, in multimodal AD, relying only on the OCBE module to get the student input results in underutilization of the rich multimodal information. In particular, when one modality is anomalous and the other is normal, we believe that using the information from the normal modality aids in reconstructing the anomaly-free features of the anomalous modality. Therefore, we propose Crossmodal Filter to optimize the OCBE process by incorporating the normal information from the other modality.
Fig. 4 depicts the structure and information flow of Crossmodal Filter (discussed here with modality B assisting modality A as an example). Our proposed Crossmodal Filter module consists of two parts: Bottleneck Projection (BP) and a modified version of the OCBE module. First, BP module aligns the teacher encoder features of modality B to the teacher features of modality A, preventing interference in the decoder process of modality A during later feature fusion. Then, the aligned modality B features are concatenated with the original teacher features of modality A. The OCBE module from RD is modified to retain its feature compression capability while expanding the input feature dimensions to accommodate the fused features.
It is worth noting that to ensure the features of modality B used for assistance are normal, we mimic the concept of OCBE and design Bottleneck Projection as a two-step process of compression and restoration. First, the three encoder features ${F^{B}_{E}}^{i},i=\{1,2,3\}$ of modality B are downsampled to $\times 8$ 8 by convolution. Then, these features are upsampled back to their original size using deconvolution to obtain ${F^{B}_{BP}}^{i},i=\{1,2,3\}$. During training, we minimize the distance between $F^{B}_{BP}$ and the teacher encoder features of modality A $F^{A}_{E}$, mapping modality B’s features into modality A’s feature space. During inference, the feature compact process of BP helps reduce the interference of anomalies in modality B, ensuring that normal information from modality B is effectively transferred to the modality A’s branch.
Figure 5: Illustration of the Crossmodal Amplifier within the branch of Modality A. The module is primarily composed of an Inverted Bottleneck Projection. Encoder features ${F_{E}^{B}}^{i},i=\{1,2,3\}$ from the other modality (Modality B) are first processed by this module to generate anomaly-aware features. These features are designed to resemble the encoder features of Modality A ${F_{E}^{A}}^{i},i=\{1,2,3\}$ in normal regions of Modality B, while exhibiting discrepancies in anomalous regions. This contrast enables the anomaly information from Modality B to be reflected in the Modality A branch. Finally, the amplified features are added to the student decoder output of Modality A to produce the final outputs ${F_{CA}^{B}}^{i},i=\{1,2,3\}$.
IV-B2 Crossmodal Amplifier
When only one modality has anomalies, multimodal fusion may lead to anomaly smoothing, resulting in false negatives, as discussed in Section IV-A. We address this issue with Multi-branch Distillation. However, fusion of the anomaly maps still cause the anomaly values to become less noticeable after summation. Our idea is that multimodal information interaction helps resolve this issue. Specifically, when modality A is normal and modality B is anomalous, if the anomaly information from modality B is incorporated into modality A’s branch, the anomaly map generated by modality A is also able to show the anomalies in modality B. This allows the anomaly in modality B to be emphasized in modality A’s branch, thus preventing the anomaly smoothing issue during anomaly map fusion.
As illustrated in Fig. 5, we design Crossmodal Amplifier. First, Inverted Bottleneck Projection (IBP) is proposed to map modality B’s features to modality A’s feature space, yielding ${F^{B}_{IBP}}^{i},i=\{1,2,3\}$. Then, the mapped features are dynamically fused with the features ${F^{A}_{D}}^{i},i=\{1,2,3\}$ which are reconstructed by the student decoder in modality A’s branch, to obtain ${F^{A}_{CA}}^{i},i=\{1,2,3\}$ by
${F^{A}_{CA}}^{i}=\frac{\exp({w_{1}^{A}}^{i})*{F^{A}_{D}}^{i}+\exp({w_{2}^{A}}^{i})*{F^{B}_{IBP}}^{i}}{\exp({w_{1}^{A}}^{i})+\exp({w_{2}^{A}}^{i})}$ (3)
where $w_{1}^{A}$ and $w_{2}^{A}$ are trainable parameters, initially set to 1.0 during training.
Inverted Bottleneck Projection only learns to map modality B’s normal features to modality A’s feature space during training. It means that IBP is not able to make the correct feature mapping when modality B has anomalies in the test sample. Therefore, the fused feature ${F^{A}_{CA}}$ includes modality B’s anomalous information, which leads to a larger distance between the teacher encoder’s features $F^{A}_{E}$ and ${F^{A}_{CA}}$, allowing the anomaly in modality B to be detected in modality A’s branch.
In contrast to Bottleneck Projection in CF, the proposed Inverted Bottleneck Projection first expands the dimensions of features and then compresses them to the original sizes. This process makes the IBP module more sensitive to distinct features, thereby better preserving the influence of anomaly information during feature mapping.
IV-C Training Objectives
The overall optimization objective of our method is composed of three main components. Firstly, similar to the approach in RD, we optimize the features output by the student decoder using cosine similarity as
$\mathcal{L}^{2D/3D}_{D}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{D}^{2D/3D}}^{i})\}$ (4)
where $\mathcal{L}^{2D}_{D}$ and $\mathcal{L}^{3D}_{D}$ are the losses corresponding to 2D and 3D branches respectively.
Secondly, to ensure the effectiveness of Crossmodal Filter, it is crucial to align the mapping features $F_{BP}^{3D/2D}$ output by Crossmodal Filters to the encoder features, as shown in Fig. 4. The training losses of 2D and 3D branches are calculated as
$\mathcal{L}^{2D/3D}_{CF}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{BP}^{3D/2D}}^{i})\}$ (5)
Finally, Crossmodal Amplifier needs to be optimized, which includes the optimization of feature mapping and feature fusion parameters $\{w_{1},w_{2}\}$. To guide this process, as in Fig. 5, we define a total loss for Crossmodal Amplifier composed of two parts: the feature mapping loss $\mathcal{L}^{2D/3D}_{IBP}$, which aligns the mapped crossmodal features with the target encoder features, and the output consistency loss $\mathcal{L}^{2D/3D}_{output}$, which is used to adjust feature fusion parameters. The total loss function is
$\mathcal{L}^{2D/3D}_{CA}=\mathcal{L}^{2D/3D}_{IBP}+\mathcal{L}^{2D/3D}_{output}$ (6)
$\mathcal{L}^{2D/3D}_{IBP}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{IBP}^{3D/2D}}^{i})\}$ (7)
$\mathcal{L}^{2D/3D}_{output}=\sum_{i=1}^{3}\{1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i})\}$ (8)
The overall loss $\mathcal{L}_{TRD}$ is expressed as
$\mathcal{L}_{TRD}=(\mathcal{L}^{2D}_{D}+\mathcal{L}^{2D}_{CF}+\mathcal{L}^{2D}_{CA})+(\mathcal{L}^{3D}_{D}+\mathcal{L}^{3D}_{CF}+\mathcal{L}^{3D}_{CA})$ (9)
Since each part has independent optimizers and inputs, the hyper-parameters are set to 1.0.
Figure 6: Inference process of our proposed method. The feature evolution visualization highlights that anomalies in each modality are effectively captured. In resulting anomaly maps, darker gray regions indicate higher anomaly scores.
TABLE I: Anomaly Detection and Localization Results (%) on MVTec 3D-AD with the Best in Bold and the Second-best Underlined
Metrics Method Bagel Cable Gland Carrot Cookie Dowel Foam Peach Potato Rope Tire Average
I-AUC BTF [7] 91.8 74.8 96.7 88.3 93.2 58.2 89.6 91.2 92.1 88.6 86.5
M3DM [8] 99.4 90.9 97.2 97.6 96.0 94.2 97.3 89.9 97.2 85.0 94.5
Shape-guided [32] 98.6 89.4 98.3 99.1 97.6 85.7 99.0 96.5 96.0 86.9 94.7
AST [5] 98.3 87.3 97.6 97.1 93.2 88.5 97.4 98.1 100 79.7 93.7
MMRD [9] 99.9 94.3 96.4 94.3 99.2 91.2 94.9 90.1 99.4 90.1 95.0
CFM [6] 98.8 87.5 98.4 99.2 99.7 92.4 96.4 94.9 97.9 95.0 96.0
TRD (Ours) 99.6 96.1 98.2 98.5 99.7 90.1 99.2 97.6 100 83.3 96.2
P-AUC BTF [7] 99.6 99.2 99.7 99.4 98.1 97.4 99.6 99.8 99.4 99.5 99.2
M3DM [8] 99.5 99.3 99.7 98.5 98.5 98.4 99.6 99.4 99.7 99.6 99.2
AST [5] - - - - - - - - - - 97.6
MMRD [9] - - - - - - - - - - 99.2
CFM [6] 99.7 99.2 99.9 97.2 98.7 99.3 99.8 99.9 99.8 99.8 99.3
TRD (Ours) 99.3 99.3 99.8 99.3 99.7 97.4 99.8 99.9 99.7 99.5 99.4
PRO BTF [7] 97.6 96.9 97.9 97.3 93.3 88.8 97.5 98.1 95.0 97.1 95.9
M3DM [8] 97.0 97.1 97.9 95.0 94.1 93.2 97.7 97.1 97.1 97.5 96.4
Shape-guided [32] 98.1 97.3 98.2 97.1 96.2 97.8 98.1 98.3 97.4 97.5 97.6
AST [5] 97.0 94.7 98.1 93.9 91.3 90.6 97.9 98.2 88.9 94.0 94.4
MMRD [9] 98.6 99.0 99.1 95.1 99.0 90.1 99.0 99.0 98.7 98.2 97.6
CFM [6] 98.0 96.6 98.2 94.7 95.9 96.7 98.2 98.3 97.6 98.2 97.2
TRD (Ours) 97.7 98.1 99.1 98.2 98.7 91.0 99.4 99.5 98.2 97.4 97.7
IV-D Inference
IV-D1 Anomaly Scoring
Following previous KD-based methods, we generate the anomaly map for each branch using cosine distance as
$\mathcal{M}^{2D/3D}=\sum_{1}^{3}(1-\mathrm{Sim}({F_{E}^{2D/3D}}^{i},{F_{CA}^{2D/3D}}^{i}))$ (10)
For the fusion of anomaly maps of two branches, we refer to the method in [2]: First, we calculate the means and standard deviations of the values in the anomaly maps using all normal samples from the validation set, denoted as $\mu^{2D/3D}$ and $\sigma^{2D/3D}$, respectively. Then, during inference, we normalize the anomaly maps for both branches as
$\mathcal{M}^{2D/3D}_{norm}=\frac{\mathcal{M}^{2D/3D}-\mu^{2D/3D}}{\sigma^{2D/3D}}$ (11)
Finally, we rescale the maps to the same scale and sum the anomaly maps from both branches to obtain the final anomaly map as $\mathcal{M}=\mathcal{M}^{2D}_{norm}+\mathcal{M}^{3D}_{norm}$. And we simply choose $\mathcal{S}=\max(\mathcal{M})$ as the anomaly score.
IV-D2 Feature Evolution Analysis
To better understand the behavior of our TRD during inference, Fig. 6 visualizes the feature evolution process within TRD. Intuitively, Crossmodal Filter effectively supplements information from the other modality while suppressing anomaly signals that may interfere with anomaly-free feature reconstruction. In contrast, Crossmodal Amplifier allows anomaly information from one modality to be reflected in the other branch. Through the joint modulation of both components, each branch becomes aware of all modality-specific anomalies, resulting in more accurate and robust anomaly detection.
V Experiments
V-A Experimental Settings
Dataset
We conduct experiments on datasets MVTec 3D-AD [4] and Eyecandies [30]. MVTec 3D-AD is a 3D industrial anomaly detection dataset featuring 10 categories of industrial objects. The training set consists of 2,656 normal samples, each including an RGB image and corresponding 3D scan information, while the validation set contains 294 normal samples. The test set comprises 1,197 samples, including 948 anomalous samples and 249 normal samples, with pixel-level annotations provided for evaluation. For our experiments, we use the processed depth maps derived from the 3D scans as the 3D modality. Eyecandies is a synthetic dataset of 10 candy categories in an industrial conveyor scenario, providing RGB images, depth maps, and surface normal maps. It contains 10k training samples, 1k validation samples, and 4k test samples. In our experiments, we select normal maps as the 3D modality and chose images under consistent lighting conditions as the 2D modality.
Evaluation Metrics
The evaluation metrics for anomaly detection include image-level area under the receiver operating characteristic curve (I-AUC) and average precision (I-AP). For anomaly localization, pixel-level AUC (P-AUC) and AP (P-AP) are used. We also employ the per-region overlap (PRO) metric calculated according to the official code [4] for anomaly localization, which treats anomaly regions of different size equally.
Implementation Details
All RGB images, depth maps and normal maps are scaled to $\times 256$ 256. The training is performed on a single Nvidia GTX 3090 GPU, and a separate detection model is trained for each category. The network backbone is WideResNet-50 [38]. The batch size is 16, and the training lasts for 200 epochs. Each branch is optimized using a separate Adam optimizer with a learning rate of 0.005. In inference, the anomaly maps of both branches are smoothed using a Gaussian filter with $\sigma=4$.
TABLE II: Anomaly Detection and Localization Results (%) on Eyecandies with the Best in Bold and the Second-best Underlined
Metrics Methods
Candy
Cane
Chocolate
Cookie
Chocolate
Praline
Confetto
Gummy
Bear
Hazelnut
Truffle
Licorice
Sandwich
Lollipop
Marsh-
mallow
Peppermint
Candy
Average
I-AUC M3DM [8] 62.4 95.8 95.8 100 88.6 75.8 94.9 83.6 100 100 89.7
AST [5] 57.4 74.7 74.7 88.9 59.6 61.7 81.6 84.1 98.7 98.7 78.0
MMRD [9] 85.4 100 94.6 99.8 90.8 74.7 96.6 98.4 100 100 94.0
CFM [6] 68.0 93.1 95.2 88.0 86.5 78.2 91.7 84.0 99.8 96.2 88.1
TRD (Ours) 86.7 100 98.1 98.1 87.2 72.8 97.1 93.0 100 99.5 93.3
P-AUC M3DM [8] 97.4 98.7 96.2 99.8 96.6 94.1 97.3 98.4 99.6 98.5 97.7
AST [5] 76.3 96.0 91.1 96.9 78.8 83.7 91.8 92.4 98.3 96.8 90.2
MMRD [9] - - - - - - - - - - 98.3
CFM [6] 98.3 98.2 96.4 98.9 94.9 94.6 96.9 98.0 99.5 98.7 97.4
TRD (Ours) 99.5 99.0 99.0 99.5 95.7 94.6 99.6 99.0 99.7 99.5 98.5
PRO M3DM [8] 90.6 92.3 80.3 98.3 85.5 68.8 88.0 90.6 96.6 95.5 88.2
AST [5] 51.4 83.5 71.4 90.5 58.7 59.0 73.6 76.9 91.8 87.8 74.4
MMRD [9] 97.5 97.0 94.2 98.5 91.7 68.0 97.0 94.1 99.0 99.2 93.6
CFM [6] 94.2 90.2 83.1 88.0/96.5 87.5 76.2 79.1 91.3 93.9 94.9 88.7
TRD (Ours) 98.0 94.1 96.8 98.1 93.6 77.3 98.1 94.4 98.9 98.1 94.7
V-B Main Results
V-B1 Results on MVTec-3D AD
Table I shows the quantitative image-level anomaly detection results and pixel-level anomaly localization results on the MVTec 3D-AD dataset. We compare our method TRD with several state-of-the-art 3D anomaly detection methods.
TRD consistently outperforms all other methods in both I-AUC and PRO metrics, achieving an average I-AUC of 96.2% and an average PRO of 97.7%. Notably, TRD demonstrates superior performance in detecting and localizing anomalies across a wide range of categories, including some challenging categories such as Cookie and Peach. A subset of qualitative results on MVTec 3D-AD is provided in Fig. 1, showing that our proposed TRD accurately localizes the anomaly regions.
Figure 7: Performance comparison of feature learning-based methods via radar chart.
TABLE III: Comparison of Feature Learning-based Methods on MVTec 3D-AD and Eyecandies datasets under the Same Hardware Setup (* indicates that MMRD has no reproducible code)
Method MVTec 3D-AD (%) Eyecandies (%)
Avg.
Frame
Rate (fps)
I-AUC P-AUC PRO I-AUC P-AUC PRO
AST [5] 93.7 97.6 94.4 78.0 90.2 74.4 88.1 13.9
MMRD* [9] 95.0 99.2 97.6 94.0 98.3 93.6 96.3 -
CFM [6] 96.0 99.3 97.2 88.1 97.4 88.7 94.5 12.0
TRD (Ours) 96.2 99.4 97.7 93.3 98.5 94.7 96.6 21.7
V-B2 Results on Eyecandies
We evaluate our method on the Eyecandies dataset, as shown in Table II. Our method TRD significantly outperforms all non-KD-based multimodal anomaly detection methods. While our TRD slightly lags behind MMRD [9] which uses both depth maps and surface normal maps as 3D modalities in anomaly detection metrics I-AUC, it consistently achieves superior performance 94.7% in anomaly localization metrics PRO.
V-B3 Comparison with Feature Learning-based Methods
We further compare TRD with feature learning-based methods including AST[5], MMRD[9], and CFM [6] on MVTec 3D-AD and Eyecandies. As in Fig. 7 and Table 7, TRD achieves the best average performance 96.6% and the highest frame rate 21.7 fps. Despite the multi-branch design, TRD remains lightweight due to its CNN-based architecture, outperforming other feature learning-based methods in efficiency.
TABLE IV: Anomaly Detection and Localization Results (%) of Different Modality Branches on MVTec 3D-AD
Branch I-AUC P-AUC P-PRO I-AP P-AP Avg.
2D 86.8 98.9 95.4 95.8 37.7 82.9
2D+CT 88.9 99.0 95.8 96.5 39.4 83.9
3D 87.1 96.2 88.8 95.2 33.7 80.2
3D+CT 87.3 97.1 90.1 95.7 34.1 80.9
2D+3D 93.9 99.3 97.4 98.0 45.2 86.8
2D+3D+CT (Ours) 96.2 99.4 97.7 98.8 45.5 87.5
Figure 8: Visualization of anomaly maps of different modality branches on MVTec 3D-AD.
Figure 9: Visualization of anomaly maps of different modality branches on Eyecandies.
V-B4 Singlemodal vs. Multimodal
We investigate the effectiveness of combining different modalities for AD through both quantitative and qualitative comparisons. We first compare singlemodal baselines (2D and 3D), their enhanced versions with Crossmodal Tuners (2D+CT and 3D+CT), and the full multimodal settings (2D+3D and 2D+3D+CT) in Table IV. Results on the MVTec 3D-AD dataset show that combining both modalities significantly improves performance. The I-AUC metric of 93.9% for using multiple modalities is higher than 86.8% for 2D only and 87.1% for 3D only, highlighting the complementary benefits of modality fusion.
To provide further insight, we visualize the anomaly maps of each branch under multimodal setup on MVTec 3D-AD and Eyecandies, as in Fig. 8 and Fig. 9. Some anomalies are hard to perceive in one modality but clearly visible in the other. Relying on a single modality may lead to missed detections, while our TRD designed for multimodal AD helps mitigate this issue through complementary information across modalities.
V-C Ablation Studies
TABLE V: Ablation Study Results (%) of Crossmodal Tuners on MVTec 3D-AD
CF CA I-AUC P-AUC P-PRO I-AP P-AP Avg.
- - 93.9 99.3 97.4 98.0 45.2 86.8
- 94.2 99.3 97.5 98.2 45.8 87.0
- 95.6 99.4 97.7 98.6 45.5 87.4
96.2 99.4 97.7 98.8 45.5 87.5
Figure 10: Qualitative ablation results of Crossmodal Tuners on MVTec 3D-AD.
V-C1 Study on Crossmodal Tuners
We perform an ablation study to evaluate the impact of Crossmodal Tuners on anomaly detection performance, as shown in Table V. The results reveal that both CF and CA contribute to improved performance, with the combination of both achieving the best results. Specifically, adding CF to the baseline improves the I-AUC from 93.9% to 94.2%, and incorporating CA also boosts the performance, achieving an I-AUC of 95.6%. The highest performance is achieved when both CF and CA are used together, resulting in an I-AUC of 96.2%.
We also provide visualizations of anomaly maps before and after incorporating Crossmodal Tuners, as in Fig. 10. Without Crossmodal Tuners, anomalies that are less prominent in a specific modality goes undetected in the corresponding branch. However, after introducing Crossmodal Tuners, this issue is significantly mitigated, resulting in a noticeable reduction in the anomaly miss rate across all modality branches.
TABLE VI: Ablation Study Results (%) of Different Downsampling Sizes in Bottlneck Projection on MVTec 3D-AD
Downsampling Size I-AUC P-AUC P-PRO I-AP P-AP Avg.
Identity 93.9 99.3 97.4 98.0 45.8 86.9
$\times 16$ 16 93.6 99.3 97.4 97.9 45.2 86.7
$\times 8$ 8 94.2 99.3 97.5 98.2 45.8 87.0
$\times 4$ 4 93.8 99.3 97.4 98.0 45.3 86.8
V-C2 Study on Bottlneck Projection in Crossmodal Filter
We investigate the effect of different downsampling sizes on the performance of the Crossmodal Filter, as shown in Table VI. When Bottleneck Projection reduces the feature size to $\times 8$ 8, the model achieves the best performance across multiple metrics. It is observed that excessively large downsampling size (e.g., $\times 4$ 4) leads to information loss, while small downsampling size (e.g., $\times 16$ 16) makes it difficult to filter out abnormal interference effectively. Therefore, we conclude that selecting an appropriate downsampling size is crucial for maintaining a balance between information preservation and interference removal.
TABLE VII: Ablation Study Results (%) of Different Channel Expansion Sizes in Inverted Bottlneck Projection on MVTec 3D-AD
Channel Expansion I-AUC P-AUC P-PRO I-AP P-AP Avg.
Identity 95.7 99.3 97.7 98.7 45.1 87.3
2×2$\times$2 × 95.6 99.4 97.7 98.6 45.5 87.4
4×4$\times$4 × 95.7 99.4 97.7 98.6 45.0 87.3
V-C3 Study on Inverted Bottlneck Projection in Crossmodal Amplifier
As recorded in Table VII, we analyze the impact of different channel expansion sizes on the performance of the Crossmodal Amplifier. The results demonstrate that a 2 times channel expansion strikes the best balance between detection performance and computational efficiency. This configuration achieves a high PRO of 97.7% and a P-AP of 45.5%, while maintaining relatively low computational overhead compared to the 4 times expansion.
TABLE VIII: Ablation Study Results (%) of Anomaly Map Aggregation Strategy on MVTec 3D-AD
Aggregation I-AUC P-AUC P-PRO I-AP P-AP Avg.
$\mathcal{M}^{RGB}_{norm}+\mathcal{M}^{Dep}_{norm}$ 96.2 99.4 97.7 98.8 45.5 87.5
$\mathcal{M}^{RGB}+\mathcal{M}^{Dep}$ 95.6 99.1 97.6 98.6 44.4 87.1
$\mathcal{M}^{RGB}\cdot\mathcal{M}^{Dep}$ [6] 94.3 99.3 97.2 98.2 43.5 86.5
V-C4 Study on Anomaly Scoring
Table VIII demonstrates that normalizing the anomaly maps from each modality before summing them performs better than directly adding or multiplying the maps. Normalization ensures that anomaly maps from different modalities are on the same scale, making fusion more effective.
Figure 11: Visualization of the features output by CF and CA.
V-C5 Visualization of Mapped Features
We visualizes the output features of CF and CA using the visualization decoders [13] in Fig. 11, which intuitively demonstrates the roles of CA and CF. Regardless of whether the initial modality is normal or anomalous, CF consistently produces normal outputs, enabling the student decoder of the other modality to generate anomaly-free results. When the initial modality is anomalous, CA introduces discrepancies with the other modality, amplifying anomalies for enhanced detection.
VI Conclusion
In this paper, Tuned Reverse Distillation is proposed, which offers a novel solution based on Knowledge Distillation for unsupervised multimodal AD task. TRD uses a multi-branch design to independently process each modality, enabling more accurate anomaly detection in each branch. Additionally, two crossmodal assistants are introduced: Crossmodal Filter helps the student decoder generate normal features, and Crossmodal Amplifier aims to amplify anomalies of one modality to branches of other modalities. Experimental results show that TRD achieves SOTA performance in both anomaly detection and localization.